meganmoney543:

2x + y = -2 x + y = 5 The y-coordinate of the solution to the system shown is _____.

meganmoney543:

help me

kasyforlife:

Is it for both y?

meganmoney543:

yes

kasyforlife:

Alr

Shadow:

They want y, so just isolate x in one equation, input it into the other equation, then solve for y. \[x + y= 5\] Subtract y from both sides \[x = 5 - y\] Input x into other equation: \[2x + y = -2\] Becomes, \[2(5 - y) + y = -2\] Distribute 2 to 5 and -y \[10 - 2y + y = -2\] Like terms (y) \[10 - y = -2\] Subtract 10 from both sides \[-y = -12\] Multiply both sides by -1 \[y = 12\]

meganmoney543:

thxx u too

Shadow:

You're welcome